# Metastable Cosmological Constant and Gravitational Bubbles: Ultra-Late-Time Transitions in Modified Gravity

### v1

@@ Abstract
The observed cosmological constant may originate as the minimum value (vacuum) $U_{min}$ of a scalar field potential, where the scalar field is frozen due to a large mass $m$. If this vacuum is metastable, it may decay to a true vacuum either at present or in the future. Assuming its decay rate $\Gamma$ is comparable to the Hubble expansion rate $H_{0}$, we estimate the scale of true vacuum bubbles and analyze their evolution. We find that their initial formation scale is sub-millimeter, and their tension causes rapid collapse if $m\gtrsim 1.7\cdot 10^{-3}\,eV$. For smaller masses, the bubbles expand at the speed of light. We extend our analysis to scalar-tensor theories with non-minimal coupling, finding that the nucleation scale of gravitational constant bubbles remains consistent with the sub-millimeter regime of General Relativity (GR). The critical mass scale remains around $10^{-3}\,eV$, similar to the minimally coupled case. A theoretical estimate at redshift $z_{obs}\sim 0.01$ suggests an observable bubble radius of $\sim 50$ Mpc, implying a gravitational transition triggered $\sim 300$ Myr ago, with a present-day size approaching $100$ Mpc. Additionally, we explore mass ranges ($m<10^{-3}\,eV$) and coupling parameter $\xi$ ranges ($10^{-8}\,eV^{2-n}-10^{-1}\,eV^{2-n}$) that lead to a variation in the gravitational constant $\Delta G/G_{N}$ within the $1\%-7\%$ range. This analysis is conducted in the framework of non-minimal coupling theories, where the coupling function takes the form $F(\phi)=1/\kappa-\xi\phi^{n}$, with $\kappa=8\pi G_{N}$ and $2\leq n\leq 9$. Finally, we review various local physics or/and transition based proposed solutions to the Hubble tension, including ultra-late-time transitional models ($z\sim 0.01$), screened fifth-force mechanisms, and the $\Lambda_{s}$ model, which features a transition at $z\sim 2$. We discuss observational hints supporting these scenarios and the theoretical challenges they face.
@@ I Introduction
In a special scenario of quintessence on the phenomenological front, the Universe is currently expanding due to a long-lasting (quintessence) false vacuum state, which may be a holdover from an earlier inflationary period [62][63]. In this regard, Landim and Abdalla have put out an intriguing scenario called metastable dark energy (MDE) [64]. The quantum tunneling from an unstable false to a stable true vacuum state is the basis of the model. As opposed to early times, as it occurs in the old inflationary model first proposed by Guth [65][66], the metastable state decay process occurs in the current low energy Universe. Due to the low temperature of the vacuum-matter phase, this transition from false to true vacuum can be discussed in the context of the semi-classical approach developed by Coleman and his collaborators [67][68].
Some advantages of the MDE model is its consistency with most data as it behaves like $\Lambda$CDM and its connection with inflation by assuming that the present false vacuum energy density is a leftover from an early inflationary epoch. Also, it is motivated by the string landscape [69]. Moreover, such a model has the potential to resolve the Hubble tension (an ultra-late transition event has been suggested as a solution) [70].
@@
    to a lower value induced by Early Dark Energy (EDE), modified gravity or dark radiation[73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90]. Thus, these models exploit the degeneracy of $r_{s}$ with $H_{0}$ in the context of the observable angular scale of the CMB sound horizon $r_{s}$
@@
    The main problem of this class is that despite fine-tuning of theoretical models that support them, they are only able to decrease the statistical significance of the Hubble tension and they cannot fully eliminate it. In addition, they favor a higher value of the matter density parameter $\Omega_{0m}$ thus worsening the $S_{8}$ tension[91, 92, 93, 94].
@@
    The main problem of this class of models is that $E(z)\equiv H(z)/H_{0}$ deformations are highly constrained by BAO and SnIa data in redshifts larger than $z\simeq 0.1$ [91, 106, 107, 20, 108, 109, 110, 111, 112, 113]. Thus, this class of models has been put in disfavor during the last few years. The exception constitutes the $\Lambda$sCDM model which shifts the deformation to high redshifts ($z\simeq 2$) and makes it abrupt so that it has a minimal effect on late data while maintaining consistency with the distance to the CMB sound horizon at last scattering [114, 115].
@@ VIII.2 Alternative Solutions to the Hubble Tension
One of the most compelling alternatives to the standard cosmological model is the $\Lambda_{s}$ model [242, 243, 115, 244, 114, 245, 246, 247, 248, 249], which introduces a sign-switching cosmological constant, $\Lambda_{s}(z)$, undergoing a rapid anti–de Sitter (AdS) to de Sitter (dS) transition at a characteristic redshift $z^{\dagger}\sim 2$ [245]. This transition effectively modifies the expansion history of the Universe post-recombination, allowing for a higher $H_{0}$ value consistent with local measurements, while simultaneously addressing the $S_{8}$ tension [247]. Unlike the traditional $\Lambda$ in $\Lambda$CDM, which remains constant and positive, $\Lambda_{s}(z)$ adopts negative values at $z>z^{\dagger}$, transitioning to positive values at lower redshifts. This behavior can be described using a hyperbolic tangent functional form, such as $\Lambda_{s}(z)=\Lambda_{s0}\tanh[\nu(z^{\dagger}-z)]/\tanh[\nu z^{\dagger}]$ [250], where $\Lambda_{s0}$ is the present-day value and $\nu$ controls the rapidity of the transition. In the limit $\nu\to\infty$, the model approaches the “abrupt” $\Lambda_{s}$ scenario, offering a minimal modification that introduces only one additional parameter compared to the standard model [245].
@@
Recent studies have provided robust theoretical underpinnings for the $\Lambda_{s}$ framework. Notably, Akarsu et al. [114, 248] propose a scalar field model with a hyperbolic tangent potential, wherein a phantom field with a negative kinetic term drives the AdS-to-dS transition. Remarkably, the step-like form of the potential prevents pathologies such as Big Rip singularities and violations of the weak energy condition (WEC) [251]. The resulting cosmological evolution remains smooth and continuous, with all key kinematical parameters ($H(z)$, $\dot{H}(z)$, and $q(z)$) evolving without discontinuities. The total energy density stays positive, and the effective equation of state remains above $-1$, ensuring physical viability [248]. While the energy density of the phantom field crosses zero, introducing a safe singularity in its EoS, the model successfully reconciles the Planck CMB and SH0ES local $H_{0}$ determinations, addressing the $H_{0}$ tension without introducing inconsistencies with large-scale structure observations [252, 34].
@@
Furthermore, extended versions of the $\Lambda_{s}$ model incorporate variations in the effective number of neutrino species ($N_{\text{eff}}$) and the total neutrino mass ($\sum m_{\nu}$) [250]. These extensions aim to examine whether late-time physics, specifically the mirror AdS-to-dS transition, suffices to alleviate cosmological tensions without necessitating deviations from the standard model of particle physics [253]. Analysis reveals that for datasets including Planck, BAO, and SH0ES, the $\Lambda_{s}$+$N_{\text{eff}}$+$\sum m_{\nu}$ model predicts $H_{0}\approx 73$ km s-1 Mpc-1, in excellent agreement with local measurements. Importantly, these results are achieved without requiring significant deviations from $N_{\text{eff}}=3.044$, maintaining compatibility with the standard cosmological model. The total neutrino mass upper bounds ($\sum m_{\nu}\lesssim 0.50$ eV) also align with neutrino oscillation experiments, reinforcing the model’s physical plausibility. Consequently, the $\Lambda_{s}$ framework emerges as a promising late-time modification to $\Lambda$CDM, capable of addressing major cosmological tensions through minimal yet impactful modifications to the Universe’s expansion history.
@@
On the other hand, the $\Lambda_{s}$ framework provides a fundamentally distinct approach, modifying the cosmic expansion history itself through a late-time anti-de Sitter to de Sitter transition. By allowing the cosmological constant to undergo a sign-switching transition at redshift $z^{\dagger}\sim 2$, this model not only resolves the $H_{0}$ tension but also mitigates the $S_{8}$ discrepancy, making it one of the few proposals capable of addressing multiple cosmological anomalies simultaneously. Theoretical developments have reinforced the physical plausibility of $\Lambda_{s}$, particularly through scalar field realizations that maintain continuity in cosmic evolution while avoiding singularities. Moreover, the model remains consistent with neutrino physics, large-scale structure observations, and early-time cosmology, positioning it as a compelling alternative to $\Lambda$CDM. Future high-precision cosmological surveys, along with potential constraints from gravitational wave standard sirens, will be instrumental in further testing the predictions of $\Lambda_{s}$ and its variants.

### v2

@@ Abstract
The observed cosmological constant may originate as the minimum value $U_{min}$ of a scalar field potential, where the scalar field is frozen due to a large mass $m$. If this vacuum is metastable, it may decay to a true vacuum either at present or in the future. Assuming its decay rate $\Gamma$ is comparable to the Hubble expansion rate $H_{0}$, we estimate the scale of true vacuum bubbles and analyze their evolution. We find that their initial formation scale is sub-millimeter, and their tension causes rapid collapse if $m\gtrsim 1.7\cdot 10^{-3}\,eV$. For smaller masses, the bubbles expand at the speed of light. We extend our analysis to scalar-tensor theories with non-minimal coupling, finding that the nucleation scale of gravitational constant bubbles remains consistent with the sub-millimeter regime of General Relativity (GR). The critical mass scale remains around $10^{-3}\,eV$, similar to the minimally coupled case. A theoretical estimate at redshift $z_{obs}\sim 0.01$ suggests an observable bubble radius of $\sim 50$ Mpc, implying a gravitational transition triggered $\sim 300$ Myr ago, with a present-day size approaching $100$ Mpc. Additionally, we explore mass ranges ($m<10^{-3}\,eV$) and coupling parameter $\xi$ ranges ($10^{-8}\,eV^{2-n}-10^{-1}\,eV^{2-n}$) that lead to a variation in the gravitational constant $\Delta G/G_{N}$ within the $1\%-7\%$ range. This analysis is conducted in the framework of non-minimal coupling theories, where the coupling function takes the form $F(\phi)=1/\kappa-\xi\phi^{n}$, with $\kappa=8\pi G_{N}$ and $2\leq n\leq 9$. Finally, we review various local physics or/and transition based proposed solutions to the Hubble tension, including ultra-late-time transitional models ($z\sim 0.01$), screened fifth-force mechanisms, and the $\Lambda_{\rm s}$ model, which features a transition at $z\sim 2$. We discuss observational hints supporting these scenarios and the theoretical challenges they face.
@@ I Introduction
In a special scenario of quintessence on the phenomenological front, the Universe is currently expanding due to a long-lasting (quintessence) false vacuum state, which may be a holdover from an earlier inflationary period [62][63]. In this regard, Landim and Abdalla have put out an intriguing scenario called metastable dark energy (MDE) [64]. The quantum tunneling from an unstable false to a stable true vacuum state is the basis of the model. As opposed to early times, as it occurs in the old inflationary model first proposed by Guth [65][66], the metastable state decay process occurs in the current low energy Universe. Due to the low temperature of the vacuum-matter phase, this transition from false to true vacuum can be discussed in the context of the semi-classical approach developed by Coleman and his collaborators [67][68]. Interestingly, an extension of the MDE model has been explored in which the metastable vacuum undergoes a transition into dark matter, potentially producing axion-like particles as a viable dark matter candidate [69]. However, it has been shown that this process would not lead to a complete transition to a dark matter-dominated phase, even in the distant future.
Some advantages of the MDE model is its consistency with most data as it behaves like $\Lambda$CDM and its connection with inflation by assuming that the present false vacuum energy density is a leftover from an early inflationary epoch. Also, it is motivated by the string landscape [70]. Moreover, such a model has the potential to resolve the Hubble tension (an ultra-late transition event has been suggested as a solution) [71]. For a comprehensive review of various dark energy models and modified gravity theories, including $f(R)$ gravity, scalar field theories, and holographic dark energy, see [72].
@@
    to a lower value induced by Early Dark Energy (EDE), modified gravity or dark radiation[75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93]. One notable proposal is the AdS-EDE model [94], which introduces an Anti-de Sitter (AdS) phase in EDE around recombination. This mechanism has been shown to further alleviate the Hubble tension by lifting the inferred $H_{0}$ value while maintaining consistency with CMB, BAO, and SNIa constraints. Thus, these models exploit the degeneracy of $r_{s}$ with $H_{0}$ in the context of the observable angular scale of the CMB sound horizon $r_{s}$
@@
    The main problem of this class is that despite fine-tuning of theoretical models that support them, they are only able to decrease the statistical significance of the Hubble tension and they cannot fully eliminate it. In addition, they favor a higher value of the matter density parameter $\Omega_{0m}$ thus worsening the $S_{8}$ tension [95, 96, 97, 98]. However, it has been shown that the increase in $\omega_{m}=\Omega_{0m}h^{2}$ in these models is a consequence of the CMB+BAO background compatibility, while $\Omega_{0m}$ itself does not shift significantly compared to $\Lambda$CDM [99].
@@
    The main problem of this class of models is that $E(z)\equiv H(z)/H_{0}$ deformations are highly constrained by BAO and SnIa data in redshifts larger than $z\simeq 0.1$ [95, 111, 112, 20, 113, 114, 115, 116, 117, 118]. Thus, this class of models has been put in disfavor during the last few years. The exception constitutes the $\Lambda_{\rm s}$ model which shifts the deformation to high redshifts ($z\simeq 2$) and makes it abrupt so that it has a minimal effect on late data while maintaining consistency with the distance to the CMB sound horizon at last scattering [119, 120].
    An alternative approach to modifying the late-time expansion history is through extensions of Horndeski gravity that include the Galileon term $G_{3}$. Such models introduce additional interactions in the scalar field Lagrangian, which can alter the cosmic evolution at both early and late times. Specifically, certain models with non-standard kinetic terms—while maintaining no-ghost conditions—modify the expansion history significantly at $z\lesssim 2$, making them a relevant candidate for addressing late-time cosmological tensions [121].
@@
  • Dark Matter: Symmetron fifth forces have recently been discussed as possible alternatives to particle dark matter [141, 142, 143, 144, 145].
@@
Modifications to gravity, particularly through non-minimal coupling of a scalar field to gravity could be an alternative approach to resolving the cosmological tensions. A recent study [159] analyzed the fit of scalar-tensor gravity models to the latest BAO data from the DESI 2024 [160, 161] release. The authors considered a selection of models from previous works and showed that non-minimally coupled theories provide an improved fit to the new DESI measurements compared to $\Lambda$CDM. Notably, their results suggest that a larger non-minimal coupling parameter is favored by the latest BAO observations, reinforcing the role of modified gravity in addressing late-time cosmological tensions. Among the models explored, the Brans-Dicke Galileon framework was identified as the best-fitting scenario, as it can naturally produce a phantom equation of state for dark energy. These insights further motivate investigations into non-minimal couplings as a viable explanation for the observed deviations in cosmic expansion.
Furthermore, [162] highlighted the role of such non-minimal couplings in explaining the latest BAO measurements from the DESI collaboration [160, 161]. These observations suggest a dynamical dark energy evolution crossing the phantom divide, which poses theoretical challenges for standard quintessence models. By analyzing the space of Horndeski gravity theories [163], the authors identified non-minimal coupling as a crucial mechanism to achieve a stable phantom divide crossing while maintaining consistency with DESI constraints. Extending this idea, a subsequent study [164] introduced the Thawing Gravity (TG) model, characterized by a non-minimal coupling function of the form $F(\phi)=1-\xi\phi^{2}$. This model offers a compelling solution to multiple cosmological tensions, including those related to the Hubble constant ($H_{0}$), the structure growth parameter ($S_{8}$), and the DESI BAO results. Remarkably, TG provides an improved fit to observational data compared to $\Lambda$CDM and predicts a $\sim 5\%$ variation in the effective gravitational constant $G_{\rm eff}$ on cosmological scales, which could be further tested by upcoming observations.
@@ VIII.2 Alternative Solutions to the Hubble Tension
One of the most compelling alternatives to the standard cosmological model is the $\Lambda_{\rm s}$ model [259, 260, 120, 261, 119, 262, 263, 264, 265, 266], which introduces a sign-switching cosmological constant, $\Lambda_{s}(z)$, undergoing a rapid anti–de Sitter (AdS) to de Sitter (dS) transition at a characteristic redshift $z^{\dagger}\sim 2$ [262]. This transition effectively modifies the expansion history of the Universe post-recombination, allowing for a higher $H_{0}$ value consistent with local measurements, while simultaneously addressing the $S_{8}$ tension [264]. Unlike the traditional $\Lambda$ in $\Lambda$CDM, which remains constant and positive, $\Lambda_{s}(z)$ adopts negative values at $z>z^{\dagger}$, transitioning to positive values at lower redshifts. This behavior can be described using a hyperbolic tangent functional form, such as $\Lambda_{s}(z)=\Lambda_{s0}\tanh[\nu(z^{\dagger}-z)]/\tanh[\nu z^{\dagger}]$ [267], where $\Lambda_{s0}$ is the present-day value and $\nu$ controls the rapidity of the transition. In the limit $\nu\to\infty$, the model approaches the “abrupt” $\Lambda_{\rm s}$ scenario, offering a minimal modification that introduces only one additional parameter compared to the standard model [262].
@@
Recent studies have provided robust theoretical underpinnings for the $\Lambda_{\rm s}$ framework. Notably, Akarsu et al. [119, 265] propose a scalar field model with a hyperbolic tangent potential, wherein a phantom field with a negative kinetic term drives the AdS-to-dS transition. Remarkably, the step-like form of the potential prevents pathologies such as Big Rip singularities and violations of the weak energy condition (WEC) [268]. The resulting cosmological evolution remains smooth and continuous, with all key kinematical parameters ($H(z)$, $\dot{H}(z)$, and $q(z)$) evolving without discontinuities. The total energy density stays positive, and the effective equation of state remains above $-1$, ensuring physical viability [265]. While the energy density of the phantom field crosses zero, introducing a safe singularity in its EoS, the model successfully reconciles the Planck CMB and SH0ES local $H_{0}$ determinations, addressing the $H_{0}$ tension without introducing inconsistencies with large-scale structure observations [269, 34].
@@
Furthermore, extended versions of the $\Lambda_{\rm s}$ model incorporate variations in the effective number of neutrino species ($N_{\text{eff}}$) and the total neutrino mass ($\sum m_{\nu}$) [267]. These extensions aim to examine whether late-time physics, specifically the mirror AdS-to-dS transition, suffices to alleviate cosmological tensions without necessitating deviations from the standard model of particle physics [270]. Analysis reveals that for datasets including Planck, BAO, and SH0ES, the $\Lambda_{\rm s}$+$N_{\text{eff}}$+$\sum m_{\nu}$ model predicts $H_{0}\approx 73$ km s-1 Mpc-1, in excellent agreement with local measurements. Importantly, these results are achieved without requiring significant deviations from $N_{\text{eff}}=3.044$, maintaining compatibility with the standard cosmological model. The total neutrino mass upper bounds ($\sum m_{\nu}\lesssim 0.50$ eV) also align with neutrino oscillation experiments, reinforcing the model’s physical plausibility. Consequently, the $\Lambda_{\rm s}$ framework emerges as a promising late-time modification to $\Lambda$CDM, capable of addressing major cosmological tensions through minimal yet impactful modifications to the Universe’s expansion history.
@@
On the other hand, the $\Lambda_{\rm s}$ framework provides a fundamentally distinct approach, modifying the cosmic expansion history itself through a late-time anti-de Sitter to de Sitter transition. By allowing the cosmological constant to undergo a sign-switching transition at redshift $z^{\dagger}\sim 2$, this model not only resolves the $H_{0}$ tension but also mitigates the $S_{8}$ discrepancy, making it one of the few proposals capable of addressing multiple cosmological anomalies simultaneously. Theoretical developments have reinforced the physical plausibility of $\Lambda_{\rm s}$, particularly through scalar field realizations that maintain continuity in cosmic evolution while avoiding singularities. Moreover, the model remains consistent with neutrino physics, large-scale structure observations, and early-time cosmology, positioning it as a compelling alternative to $\Lambda$CDM. Future high-precision cosmological surveys, along with potential constraints from gravitational wave standard sirens, will be instrumental in further testing the predictions of $\Lambda_{\rm s}$ and its variants.